# Generative Location Modeling for Spatially Aware Object Insertion

### v1

@@ 4.2 Object Insertion
Figure 4: Quantitative evaluation on the OPA dataset, higher and to the right is best. For instruction-tuned approaches, each dot represents a different guidance scale ranging from 2 to 10. For other methods, guidance scale has a negligible effect, and we therefore show a single point.
Baselines. We compare our object insertion pipeline against recent instruction-tuned image editing models, which represent some of the most advanced and capable techniques for object insertion to date. Specifically, we evaluate against general-purpose image editing models InstructPix2Pix (Brooks et al., 2023), HIVE (Zhang et al., 2024b), and MagicBrush (Zhang et al., 2024a), as well as object insertion-specific models, Paint-by-Inpaint (Wasserman et al., 2024) cand Diffree (Zhao et al., 2024).
Moreover, we experiment with three additional explicit location models that we also pair with the PowerPaint inpainter. First, we test random locations as a simple baseline. Then, we train ContextCNN (Dvornik et al., 2018), a classifier designed to assess masked regions of an image, determining their suitability for object placement. To be able to use it with the OPA class space, we retrain the model on COCO (whose classes comprehend all the ones in OPA. Furthermore, we train a Faster-RCNN object detector (Ren et al., 2015) using positive OPA annotations to serve as a high-performing discriminative model.
Evaluation Metrics. Similar to existing image editing benchmarks (Wasserman et al., 2024), we evaluate the success of object insertion by considering both how well the original scene is preserved and how accurately the target object is inserted. Using an object detector (Zhu et al., 2021) we separate the background, which ideally should remain untouched, and the foreground, where the newly inserted object appears.
To measure background preservation, we compute the Structural Similarity Index Measure (SSIM) (Wang et al., 2004) on the background region. To assess the accuracy of the object, we measure the CLIP similarity (Radford et al., 2021) between the cropped object and the text “an {object class}”. If no new object is detected, we assign a CLIP score of 0 to reflect the failure to properly insert the object.
For the PIPE dataset, we also evaluate the diversity of edits by calculating the average LPIPS distance (Zhang et al., 2018) across 10 different edits per background-object pair. High LPIPS indicates that a model can generate diverse results from the same instruction, highlighting approaches that are limited to producing a single edit.
Results. We compare instruction-finetuned image editing models, location models paired with strong inpainting models, and our approach in Figure 4. Our approach substantially outperforms all baselines by leveraging a dedicated location model for inpainting, which effectively reduces background distortions while maintaining high-quality object generation.
@@
The instruction-tuned models often face a trade-off: they either preserve the background well but fail to generate the object convincingly, or they successfully generate the object but significantly alter the surrounding scene. This inconsistency may stem from the fact that these models address both the placement and generation tasks simultaneously, leading to suboptimal object locations. In contrast, using a dedicated location model allows one model to focus on spatial reasoning, while the inpainting method concentrates on rendering realistic objects.
Table 1: Evaluation on the PIPE benchmark.
BG SSIM FG CLIP LPIPS
@@
Figure 5: User study results on edited images (OPA). Left, blue: ours preferred. Right, red: baseline preferred. Middle, grey: no preference.
@@
We further evaluate our approach on the PIPE test set (Wasserman et al., 2024), which contains images with specific objects removed. Using the same metrics reported in Table 4.2, our results are consistent with those observed in the OPA test set. Notably, Paint-by-Inpaint (Wasserman et al., 2024) lacks diversity, frequently generating identical outputs, as shown in Figure 8. This is likely due to training on datasets created by object removal through inpainting. Moreover, we appreciate how InstructPix2Pix and HIVE achieve very diverse edits. However, we notice that this result is typically achieved by generating entirely new images, rather than editing the input scene, as also testified by their low background SSIM scores.
User Study. To further validate that our model’s insertions are favored by human observers, we perform a user user study comparing our method to the four strongest baselines. We show pairs of edited OPA images and ask users to indicate which of the two is the better edit, for 46 participants, each of which ranks 40 image pairs. For further details, see Appendix Section E. Results are shown in Figure 5. Participants preferred edits generated by our approach over those from baseline approaches, indicating that our metrics agree with human preference, and that better edit quality can be achieved through precise location modeling. Additional qualitative examples can be found in Figure 7, Figure 8, and Appendix Section F.
4.3 Location Modeling
Baselines. We compare our generative location model against two discriminative approaches that classify locations as plausible or implausible, namely ContextCNN (Dvornik et al., 2018) and our Faster-RCNN (Ren et al., 2015) based location model, as described in Section 4.2. Additionally, we compare to two object placement baselines (Zhou et al., 2022; Niu et al., 2022) that perform placement of specific object segments rather than generic class labels. We use the official implementations relying on foreground segments available within the OPA dataset. Note that a direct comparison to these two methods is challenging, as our method does not use the object segment, and it is hard to say whether access to one makes the task harder or easier. Still, we include these baselines for completeness.
Figure 6: (Left) Example evaluation scenario. Predicted boxes are counted only if a positive or negative ground-truth box meets an IoU above the threshold. (Right) TPR-FPR curves. Each line is constructed by increasing the number of sampled locations from $\{10,20,\ldots,100\}$. Higher and to the left is better.
Evaluation Metrics. The OPA test set provides plausible (positive) and implausible (negative) locations for objects given an image. However, due to the sparse nature of these annotations, it is impossible to sample locations until every ground-truth bounding box is matched with a prediction. Therefore, we evaluate location models based on a “hit rate” metric, which compares the rate of each predicted box being a plausible or implausible location. Specifically, we measure the True Positive Rate (TPR) and False Positive Rate (FPR) for a given set of predicted locations. Given $K$ predicted locations, we match them to the ground-truth labels using the Hungarian algorithm (Kuhn, 1955), where the cost function is the inverse of the intersection over union (IoU).
True positive predictions are defined as predictions assigned to positive labels with an IoU above 0.7, while false positive predictions are those assigned to negative labels under the same IoU threshold. Any positive or negative ground-truth locations that are not matched are counted as false negatives and true negatives. Predictions that do not correspond to any labeled locations are ignored, as their true labels cannot be determined. We refer the reader to Figure 6 (left) for an example of such assignments. TPR and FPR are then computed using standard definitions. Intuitively, TPR represents the rate of a predicted location being a correct (positive) location, and FPR represents the rate of a predicted location being an incorrect (negative) location.
Results. We plot the TPR and FPR for different number of sampled locations $K=10,20,\dots,100$ in Figure 6 (right). Our generative location model consistently achieves a higher TPR at the same FPR, appearing in the top-left region of the plot. In contrast, discriminative baselines (Dvornik et al., 2018; Ren et al., 2015) fail to reach a high TPR, even after predicting 100 locations, possibly due to the penalization of unlabeled positive locations during training. Our generative approach avoids any assumption about unlabeled locations, allowing it to achieve higher accuracy in identifying plausible ones. This finding suggests the effectiveness of generative modeling in scenarios where the sparsity in annotations hinders training a discriminative model.
4.4 Ablation Study
DPO Training. Training a generative location model exclusively on positive locations already demonstrates strong performance, as shown in Figure 6 (right). However, incorporating negative labels further enhances accuracy by explicitly guiding the model on where not to predict object locations. Notably, unlike existing location modeling techniques, which assume non-labeled locations as negative even when negative labels are present, our approach leverages only the negative labels provided by annotators. This ensures more precise predictions, as it avoids the potential inaccuracies introduced by assuming non-labeled locations are negative.
Alternative Inpainting Methods. Our location model is not tied to a single inpainting method and can incorporate various inpainting techniques. To illustrate this, we use the same predicted locations to render objects with the inpainting model of GLIGEN (Li et al., 2023). The resulting images rendered achieve a background SSIM of 0.6511 and foreground CLIP of 0.2833, compared to a performance of 0.7184 background SSIM, 0.2849 foreground CLIP, when using PowerPaint. Obtaining similar foreground CLIP score indicates that inpainting is often successful even with the older GLIGEN model. We therefore expect that our approach benefits from future advancements in inpainting techniques as well.
@@ 4.4 Ablation Study
Figure 8: Comparison between our method + powerpaint, and instruction-guided image editing models on the PIPE dataset. Additional results are available in Section F. Best viewed electronically.
@@ 5 Discussion
Inference Time. Our location model introduces a minimal overhead to the overall object insertion process. To measure this, we compare the inference time of the location model relative to the time required for rendering the image (i.e., inpainting). On an average across 100 runs, our model takes 0.03 seconds to sample a single location on a Nvidia Tesla V100 GPU, a minor addition compared to the 7.10 seconds needed to render an image using a 50-step diffusion reverse process for a StableDiffusion v2.1 checkpoint (Rombach et al., 2021). Paying a small upfront cost for identifying a plausible location leads to a significant improvement in the quality of object insertion.
Controlled Location Sampling.
Figure 9: Controlled location sampling for positional instructions, achieved by restricting the allowed sampling region. Instruction-tuned models often fail to follow positional instructions.
A location model not only automates the process of determining where to insert objects but also allows users to specify locations that meet particular requirements. Current instruction-tuned image editing models struggle with positional instructions, such as “Add an apple to the top-left of the image.” In contrast, our location model can handle these requests by constraining model outputs to specific areas, such as by sampling only the top-left coordinates. As illustrated in Figure 9, by factorizing the object insertion process into a location modeling step and an inpainting step, our approach offers increased control over object placement.
Limitations. Although localized object insertion helps minimize distortions in the background, predicted bounding box locations may sometimes occlude more of the background than necessary. As seen in Figure 10, large bounding boxes may lead to unwanted changes in the scene, such as changes to the background or inadvertently occluding foreground objects. It may also result in unrealistic background effects, such as missing shadows or reflections. These issues could potentially be mitigated by predicting fine inpainting masks within the bounding box regions or by developing inpainting techniques tailored specifically for object insertion, which we leave as future work.
Figure 10: Example failure case observed when inpainting large areas, for instruction “add a horse”.
@@ 6 Conclusion
In this paper, we present a location model that identifies plausible locations for objects to be inserted within an image. By taking a generative approach, we are able to work with sparsely annotated location datasets. We also show that it is beneficial to use negative labels via direct preference optimization. Our experiments suggest that in the task of object insertion, separating the problems of where to place an object and what to place in the given location, is currently much more reliable than instruction-tuned insertion approaches.
More generally, we believe that building spatial awareness is a key factor for building reliable models interacting with the real world. This is true whether a model operates in the image editing setting, as in this work, or in more complex domains, such as robotics or virtual reality. Our work shows that a model is able to learn such awareness from example positive and negative annotations. Although we focus on locations for 2D bounding boxes conditioned on images, we hope to see similar models scale to handle other scene representations (e.g., depth or semantic layouts) and precise locations (e.g., 3D bounding boxes or object masks) in the future.
Broader Impact Statement
Spatial awareness is crucial for real-world applications where precise and context-aware object placement is necessary. Improving the reliability of object insertion methods can benefit numerous fields, such as augmented reality, robotics, and generative data augmentation, by enabling more realistic and practical scene manipulations. However, image editing methods can also be misused, for example, to manipulate images by inserting individuals into scenes to damage their reputation. Our method, which specifies only the desired object class rather compositing a background image and a specific target object, is hopefully less suitable for such malicious use cases. Nevertheless, one should consider malicious use cases of image editing methods before deployment, especially if used in real world or safety-critical applications.
Reproducibility Statement
We provide comprehensive details of our architecture and data preprocessing steps in Section 4.1 and Section D. Our architectures are based on open-source implementation of GPT-2 (Radford et al., 2019), for example found at this github repo, and uses standard open source vision backbones to initialize the vision and text encoders. Furthermore, all datasets used in our experiments (Liu et al., 2021; Wasserman et al., 2024) are publicly available, facilitating the replication of our work by the community.
@@ A.1 PIPE Dataset
The PIPE dataset (Wasserman et al., 2024) was created by removing objects from object detection datasets (Lin et al., 2014; Kuznetsova et al., 2020; Gupta et al., 2019) using an inpainting model, resulting in over 600 object classes for insertion and more than 888,000 training pairs of images, showing scenes before and after object removal. To preprocess this dataset, we pair each background image with the original location of the removed object. The locations are identified by computing the pixel-wise difference between the before-and-after images, and extracting coordinates where the difference exceeds a certain threshold. Despite this thresholding, the resulting bounding boxes can be noisy, and the background images generated by the inpainting model often contain artifacts. Also, the PIPE dataset only provides a single positive location for an object for each background image and does not provide negative labels.
@@ A.2 OPA Dataset
The OPA dataset was created by manually annotating samples from the COCO dataset (Lin et al., 2014), resulting in 47 object categories for object placement. The dataset is intended for the task of finding locations to copy-paste images of objects, and therefore includes background images, object images with transparent backgrounds, and labeled plausible/implausible locations for placing the objects. Since our focus is on object location modeling rather than the insertion of object images, we ignore the object images and restructure the dataset to include pairs of background images and their corresponding object locations. This restructuring yields 1,496 training samples and 184 test samples. While the number of images is relatively small, each sample contains around 40 annotations, making it a richly annotated dataset. In total, the OPA train set contains 21,376 positive labels and 40,698 negative labels.
Despite having an average 40 annotations for each sample, this accounts for fewer than 1% of the typical number of anchor boxes used in object detectors, leaving most locations unlabeled. Furthermore, as illustrated in Figure 11, not only are the number of bounding boxes across samples highly imbalanced, but the distribution of positive and negative labels for each sample is extremely inconsistent. This sparsity and imbalance make it extremely challenging to train discriminative models that classify locations as plausible or implausible. Our generative approach bypasses this issue by modeling the distribution of plausible locations, using negative labels only for DPO.
@@ Appendix B Does Location Matter?
To demonstrate that the quality of locations directly influences the quality of generated objects, we present a proof-of-concept experiment in Section 12. For each image in the OPA dataset, we uniformly sample 10 random locations and then interpolate between these random locations and manually annotated ones, performing inpainting at six distinct interpolation points. To evaluate the success of inpainting, we measure the foreground CLIP similarity (Radford et al., 2021) between the cropped object and the text “an {object class}”. For reference, we also include the performance of MagicBrush, which inserts objects without explicit location modeling. As the inpainting locations become more precise, the fidelity of generated objects increases, underscoring the importance of accurate locations. Inpainting in incorrect locations often results in failed insertions, motivating the development of a dedicated location model to provide spatial awareness for inpainting models.
@@
Appendix C Alternative Approaches for Location Modeling
Vision-language models (VLMs) can also be used to predict plausible locations for object placement based on image inputs. Although these models can generate bounding box coordinates as part of their responses, we find that they are largely ineffective at predicting meaningful locations, with their bounding box outputs being comparable to random guesses. Specifically, we use LLaVA-13B (Liu et al., 2024) using the following prompt:
USER: <image>If a new {object_name} would appear in this scene,
what would be the coordinates of the {num_samples} different
plausible locations? Answer in JSON format
@@ Appendix C Alternative Approaches for Location Modeling
ASSISTANT: In this scene, {num_samples} most plausible locations
of a newly inserted {object_name} are:
We also compare against object placement approaches (Zhou et al., 2022; Niu et al., 2022) that predict locations based on both the background image and a tightly masked image of the object. Since these models are provided with the exact aspect ratio of the object, they only need to predict the location and scale of the bounding box. As previously mentioned in Section A, the OPA dataset includes object images, allowing us to measure the TPR and FPR on the same test set. Despite having the advantage of having provided with ground-truth aspect ratios, these models are outperformed by our location model, which demonstrates superior performance even without access to object images. A full comparison of these location models are plotted in Figure 13.

### v2

@@ 4.3 Object Insertion
Baselines.
We evaluate our location model in an object insertion pipeline, where we first identify plausible insertion locations using our generative location model and subsequently render objects at these locations using an off-the-shelf inpainting model, PowerPaint [55]. Our method is evaluated against several strong baselines, including general instruction-based image editing models such as InstructPix2Pix [2], HIVE [49], and MagicBrush [46], as well as specialized object-insertion models like Paint-by-Inpaint [43] and Diffree [51]. To further isolate and demonstrate the advantages of our generative location model, we additionally experiment with Faster-RCNN [36], a discriminative location model, integrated into the same insertion pipeline.
These comparisons illustrate that explicitly modeling plausible object locations significantly enhances insertion realism and accuracy, providing a clear advantage over general-purpose editing approaches and other less accurate location models.
Evaluation Metrics.
Similar to existing image editing benchmarks [43], we evaluate the success of object insertion by considering both how well the original scene is preserved and how accurately the target object is inserted. Using an object detector [54] we separate the background, which ideally should remain untouched, and the foreground, where the newly inserted object appears.
To measure background preservation, we compute the Structural Similarity Index Measure (SSIM) [42] on the background region. To assess the accuracy of the object, we measure the CLIP similarity [34] between the cropped object and the text “an {object class}”. If no new object is detected, we assign a CLIP score of 0 to reflect the failure to properly insert the object.
For the PIPE dataset, we also evaluate the diversity of edits by calculating the average LPIPS distance [48] across 10 different edits per background-object pair. High LPIPS indicates that a model can generate diverse results from the same instruction, highlighting approaches that are limited to producing a single edit.
Results on OPA.
We compare instruction-finetuned image editing models, location models paired with strong inpainting models, and our approach in Figure 6. Our approach substantially outperforms all baselines by leveraging a dedicated location model for inpainting, which effectively reduces background distortions while maintaining high-quality object generation.
@@
The instruction-tuned models often face a trade-off. They either preserve the background well but fail to generate the object convincingly, or they successfully generate the object but significantly alter the surrounding scene. This inconsistency may stem from the fact that these models address both the placement and generation tasks simultaneously, leading to suboptimal object locations. In contrast, using a dedicated location model allows one model to focus on spatial reasoning, while the inpainting method concentrates on rendering realistic objects.
BG SSIM $\uparrow$ FG CLIP $\uparrow$ LPIPS $\uparrow$
@@
Table 1: Evaluation on PIPE. BG SSIM represents background similarity, FG CLIP measures faithful generation of the object, and LPIPS measure the diversity of image across multiple generations.
Results on PIPE.
@@
We further evaluate our approach on the PIPE test set [43], which is a dedicated dataset for evaluating object insertion methods. Using the same metrics reported in Table 1, our results are consistent with those observed in the OPA test set. Notably, Paint-by-Inpaint [43] lacks diversity, frequently generating identical outputs, as seen in extremely low LPIPS scores. This is likely due to training on datasets created by object removal through inpainting. Moreover, although we appreciate how InstructPix2Pix and HIVE achieve very diverse edits, we notice that this result is typically achieved by generating entirely new images, rather than editing the input scene, as also testified by their low background SSIM scores.
User Study.
To evaluate the quality of object placement from a human perceptual perspective, we conducted a user study comparing our method against four competitive baselines. We presented 46 participants with pairs of edited images from the OPA dataset, asking each participant to select the image with the more realistic and coherent object insertion. Each participant evaluated 40 image pairs. Further details can be found in Appendix Section F.
The results of the user study, summarized in Figure 7, clearly demonstrate that our approach is consistently preferred over the baselines. Notably, our generative location model significantly outperforms the strongest discriminative baseline, FasterRCNN [36], with twice as many participants preferring edits guided by our model’s predicted locations. These findings underscore the critical impact of accurate location modeling, highlighting that improvements at the placement stage propagate substantially into downstream applications reliant on realistic object locations.
Figure 7: User study on edited images (OPA). Blue: ours preferred. Red: baseline preferred. Grey: no preference.
@@
Ablation study.
Training a generative location model exclusively on positive locations already demonstrates strong performance, as shown in Figure 5, Ours w/o DPO. However, incorporating negative labels further enhances accuracy by explicitly guiding the model on where not to predict object locations. Notably, unlike existing techniques which assume non-labeled locations as negative even when negative labels are present, our approach leverages only the negative labels provided by annotators. This ensures more precise predictions, as it avoids the potential inaccuracies introduced by assuming non-labeled locations are negative.
@@ 5 Discussion
Inference Cost.
Our location model introduces a minimal overhead to the object insertion process. To measure it, we compare the inference time of the location model relative to the time required for rendering the image (i.e., inpainting). On an average across 100 runs, our model takes 0.03 seconds to sample a single location on a Nvidia Tesla V100 GPU, a minor addition compared to the 7.10 seconds needed to render an image using a 50-step diffusion reverse process for a StableDiffusion v2.1 [38]. Paying a small upfront cost for identifying a plausible location leads to a significant improvement in the quality of object insertion.
Limitations.
While our generative formulation enables effective training of location models with sparse annotations, it is important to note that these annotations remain costly to collect. As a result, our approach is currently limited to datasets that provide such annotations. A growing trend, involves the automatic generation of such data by removing objects from scenes (e.g. PIPE dataset). In the future, we aim to expand these datasets by incorporating multiple object removals, thus enabling a more comprehensive modeling of plausible locations across various scenarios.
@@ 6 Conclusion
In this work, we propose a generative location model designed to identify plausible locations for inserting hypothetical objects into existing scenes. Unlike previous discriminative approaches, our generative formulation naturally addresses the inherent sparsity and ambiguity of location annotations, enabling robust and accurate predictions even when labeled data is limited. Our experiments demonstrate that our model effectively captures the distribution of plausible locations. We perform additional experiments on a downstream task, object insertion. These experiments confirm that accurately predicted locations significantly enhance the realism and coherence of synthesized images, underscoring the crucial role that high-quality location modeling plays in successful object insertion.
More generally, we believe that building spatial awareness is a key factor for building reliable models interacting with the real world. This is true whether a model operates in the image editing setting, as in this work, or in more complex domains, such as robotics or virtual reality. Our work shows that a model is able to effectively learn such awareness from example positive and negative annotations, following a generative modeling of these locations. Although we focus on locations for 2D bounding boxes conditioned on images, we hope to see similar models scale to handle other scene representations (e.g., depth or semantic layouts) and precise locations (e.g., 3D bounding boxes or object masks) in the future.
@@ A.1 PIPE Dataset
The PIPE dataset [43] was created by removing objects from object detection datasets [27, 22, 12] using an inpainting model, resulting in over 600 object classes for insertion and more than 888,000 training pairs of images, showing scenes before and after object removal. To preprocess this dataset, we pair each background image with the original location of the removed object. The locations are identified by computing the pixel-wise difference between the before-and-after images, and extracting coordinates where the difference exceeds a certain threshold. Despite this thresholding, the resulting bounding boxes can be noisy, and the background images generated by the inpainting model often contain artifacts. Also, the PIPE dataset only provides a single positive location for an object for each background image and does not provide negative labels.
@@ A.2 OPA Dataset
The OPA dataset was created by manually annotating samples from the COCO dataset [27], resulting in 47 object categories for object placement. The dataset is intended for the task of finding locations to copy-paste images of objects, and therefore includes background images, object images with transparent backgrounds, and labeled plausible/implausible locations for placing the objects. Since our focus is on object location modeling rather than the insertion of object images, we ignore the object images and restructure the dataset to include pairs of background images and their corresponding object locations. This restructuring yields 1,496 training samples and 184 test samples. While the number of images is relatively small, each sample contains around 40 annotations, making it a richly annotated dataset. In total, the OPA train set contains 21,376 positive labels and 40,698 negative labels.
Despite having an average 40 annotations for each sample, this accounts for fewer than 1% of the typical number of anchor boxes used in object detectors, leaving most locations unlabeled. Furthermore, as illustrated in Fig. 9, not only are the number of bounding boxes across samples highly imbalanced, but the distribution of positive and negative labels for each sample is extremely inconsistent. This sparsity and imbalance make it extremely challenging to train discriminative models that classify locations as plausible or implausible. Our generative approach bypasses this issue by modeling the distribution of plausible locations, using negative labels only for DPO.
@@ Appendix B Does Location Matter?
To demonstrate that the quality of locations directly influences the quality of generated objects, we present a proof-of-concept experiment in Section 10. For each image in the OPA dataset, we uniformly sample 10 random locations and then interpolate between these random locations and manually annotated ones, performing inpainting at six distinct interpolation points. To evaluate the success of inpainting, we measure the foreground CLIP similarity [34] between the cropped object and the text “an {object class}”. For reference, we also include the performance of MagicBrush, which inserts objects without explicit location modeling. As the inpainting locations become more precise, the fidelity of generated objects increases, underscoring the importance of accurate locations. Inpainting in incorrect locations often results in failed insertions, motivating the development of a dedicated location model to provide spatial awareness for inpainting models.
@@
Appendix C VLMs for Location Modeling
Vision-language models (VLMs) can also be used to predict plausible locations for object placement based on image inputs. Although these models can generate bounding box coordinates as part of their responses, we find that they are largely ineffective at predicting meaningful locations, with their bounding box outputs being comparable to random guesses. Specifically, we use LLaVA-13B [28] using the following prompt:
USER: <image>If a new {object_name} would appear in this
scene, what would be the coordinates of
the {num_samples} different plausible locations?
Answer in JSON format
@@ Appendix C VLMs for Location Modeling
ASSISTANT: In this scene, {num_samples} most plausible
locations of a newly inserted {object_name} are:
We also compare against object placement approaches [52, 30] that predict locations based on both the background image and a tightly masked image of the object. Since these models are provided with the exact aspect ratio of the object, they only need to predict the location and scale of the bounding box. As previously mentioned in Section A, the OPA dataset includes object images, allowing us to measure the TPR and FPR on the same test set. Despite having the advantage of having provided with ground-truth aspect ratios, these models are outperformed by our location model, which demonstrates superior performance even without access to object images. A full comparison of these location models are plotted in Fig. 11.